Testbeam June&July 2011

Software repositories and Storage: /castor/cern.ch/user/r/rd51

Team (mpgd-beam-tech@cernNOSPAMPLEASE.ch)

Marcello Abbrescia, Eraldo Olivieri (16 2474), Stefano Colafranceschi (16 2881), Vindhya Engu, Esha Sarkar,Alfredo Gutierrez, Marcello Maggi, Andrey Marinov, Ankit Durgasis Mohapatra, Rajesh Pandey, Archana Sharma (16 4875), Michael Staib, Michal Zientek, Jasbir Singh, Leszek Ropelewski, Christopher Armaingaud, Jeremie Merlin, (and once in a while) Michael Tytgat

Other useful telephone numbers: H2 control room = 79047, CCC = 77500

Runlist (GE_11proto2 + GE_11proto1 + timingGEM)

Excel table with summary runlist

DAILY ELOGS

10/06/2011

DAY: Hardware installation finished, network system is debugged, HV and LV cables installed/tested, tracker on gas since days ready to go ON, prototypes on gas since 15.00.

EVENING/NIGHT: Added/configured one more turbo (now 3 in total) so that we can measure time res. of GE11new-old and timingGEM all together. Solved HV connectivity problem of the GE11new (hv filter bad found). Pc backup and cleaning completed. Timing DAQ: dummy run tried (long fiber optics works fine).

11/06/2011

EARLY MORNING: Finishing preparation for dummy data-taking. Mapping electronics chip, powering up LV and HV (slowly on GE11old&new). No beam before 15.00 (problem from the PS), anyway as last night we expect something around 30counts/spill..

DAY: PS has problems, rumors about no beam.. Some problem with few readout channels on the GE11new (investigating..); all big detectors are washed with 25 volumes. At 12.00 GE11new&old at 1kV, at 18:00 GE11new&old at 2kV. Working on the software side..

EVENING: 19.00 Beam is coming (parasitically) at 21.00.. still working on the software for the new configuration.

NIGHT: Beam is not coming at all to the North Area, expert is working on site but no time estimation. In the mean while work on software adaption for the prompt analysis almost finished, positive result with dummy (empty run). Noise on the GE11new is a bit higher than the one measured in the lab, investigating this working on different grounding schema.

12/06/2011

EARLY MORNING: 08.00 CMS-MPGD H2B main user, beam ok but waiting for closing area.

DAY 11:30 Area is closed, waiting for beam. Attention: All of us should follow the new access course and be authorized to open and close area. Will get ne scheduled soon and inform. For now, we have to be careful not to go in free access mode, rather entry with key mode.

14.00 After some trouble setting the beam now we see the tracker beam profile...but alignement problems..

15.26 GE11new@ 3kV

17.00 GE11new@3NOSPAMPLEASE.5kV, alignement almost ok now.. still taking few runs.

19.00 Software bug (tricky) found and solved, beam runs ok, allignement of CMS GE1/1 new&old, taking data..

20.30 Beam steering, trying to get cleaner data, playing wih thr with the tracker, we have still misalignment issues to understand, working on it.

24.00 Continuing to understand the strange beam profile..

13/06/2011

03:00 Disentangling problems.. mapping is ok (root==labview==inspection), tracking reconstruction tuned but still few good tracks found. Beam profile is weird (measured angle of 0.3-0.1 with a misalignment of 8mm over 40cm).. and tracker is noisy and not very efficient. Checking with high statistics run with as much as possible "clean" beam.

04.40 After several checks, plots, fine tuning the problem has been addressed, there is one hole in the beam!!!!!!! Any problem in the tracker is excluded now, the only problem is either an obstacle in the beam line or some effect due to the target?!? This happens with (pure) muons and/or pions..

05.30 Setting offsets for ge11new&old and timingGEM, difficult alignement..

07.00 GE1/1new&old aligned, nice beam profile, TimingGEM has been rotated 90degree (it was installed in wrong position), but it gives no hit and the chip does not show anything in thelatency scan. On the full-size prototype now sectors to be scanned are shifted to the right side (right and left are simmetric) just to make alignement with the tracker easier.

07.15 Switching OFF LV, HV, control room locked.

REMAINING TO DO.. -Measuring distance between each trackerGEM with each detector in the Magnet. -Debug/Check position of the TimingGEM which does not show up. -Once this is done we start the scan.

11.50 We are going to check the timing GEM VFAT. All the detectors are on - (HV,LV)

12.29 Moved the timing GEM VFAT 7cm down

12:34 Switched on the high voltage.

13.30 Setting the beam, target was not in correct place, looking for nice settings.. found, nice beam profile.

17.00 Waiting for the configuration file for DAQ.. in the mean while beam quality is degradated..

20.00 Files are finally ready, loading into the DAQ but working to improve beam for the night. Tracker2 increased of 50V.

22.30 Finally H2B.987 FM-H2.B-1.06mr works very nicely, beam profile and rate are good, data-taking is faster!!

23.00 Actually the rate now is fine but the TDC does not stand it.. quick access to prescale the TDCs and check the alignement.

23.30 During the access we realized a very high rate (10times bigger than with BEAM-ON).. checking this safety issue with CCC

24.00 No beam..

14/06/2011

00.15 Huge rate! It is going just fast.. now I am really busy to look runs to validate data-quality, runs are piling up..

00.50 Radiation Alarm, intensity of beam on the target too high, that's why the rate is huge, we agreed to reduce intensity beam, improve their quality and hopefully maintain high rate.

1.00 Prompt analysis: Tracker1&Tracker3 increased of 50V because same action on Tracker2 was effective.

1.30 Installing the prescaler to get TDC working with any rate, checking beam quality with reduced intensity 20.3E11 (instead of 60E11)

3.30 HVscan P1-P2-P3 done, 50% scan is performed.

7.00 Still taking data.. no problem but rate is not so high..

13.30 HV scan finished, the decision is to continue performing additional HV scan in the high and low eta part of the detector

14.30 Moving the detector to the high eta part and starting data-taking

16.00 During the meeting we agreed to finish these additional HV scan with 10k events (instead of 30k) and change gas over evening/night

18.00 Beam has problems.. CCC experts are working on this..

22:25 Beam was restored at 8 p.m. We did not see any signal from our three prototype GEMs. beam was constantly interrupted making it impossible to debug the system.

00:35 The beam is back at 13,000 events (pions) per spill. Now we are doing tests for high eta and we are on section 1. It is estimated to be completed by 6 AM. Then we will move the detectors on. So we have a meeting here at 6 AM in the morning.

15/06/2011

06:00 Due to a LV trip of the VFAT for Proto2 we have to repeat point 6 of the HV scan (last point). Then we can start the detector moving to the Low eta part. The beam is stable with nice profile.

07:00 Problem with VFATs We are replacing the bad one. The HV Scan is still not finished yet.

07:30 We have replaced the chip on Proto2 with VFAT id H0238 (it was H0264) Now we are doing tests.

07:40 The VFAT H0238 works fine. We are resuming the datataking

8:43 We have finished all the measurements at Hi ETA side. The beam is off for entire day due to the cable work. We had some problems with the prototype 2 VFAT chip (tripping) at 6th point of the measurements. We solved it by repleacing the chip.

We will use this dead time to change gas; and learn to work with the magnet; check if all equipment inside area is not affected by B field

8:59 We are now flushing ArCO2CF4 (45:15:40). Things to be done before magnet tests:

- data back up of the PC.

9:05 Problems with the CF4 botle.

10:30 Problem solved, we have gas mixture ArCO2CF4 flushing the 3 prototypes (flow: 5l/h).

12:02 There will be no magnetic field at least till tomorrow morning (waiting for an expert).

20.00 In 20 minutes there will be beam. Patrolling the area.

20:45 Switching on High Voltage.

22:30 Beam restored.

22:45 After few impressive runs (45k counts per spill), the beam has been cut (vacum problem).

23:23 Serious problem with the beam for north area (specialist called). CCC sets up the machin to extract to the other users.

23:53 Official information from CCC (confirmed many times): "No beam untill tomorrow morning for north area, more news tomorrow morning". We are switching off all the detectors and locking the counting room.

16/06/2011

8.16 CCC says no beam probably for the morning

9.00 Magnet could be ready, problem identified on some bad electrical connection.. cleaning them

10.30 Ramping up the magnet! 11.20 Vacuum problem in TDC 2-No beam to North area before 18.00 hrs-More news at afternoon

11.30 News from the SPS User Meeting: This vacuum problem will be not solved before 18.00 but it could take up to 3 days in the worst unlucky case

18.30 No beam before tomorrow, we will resume tomorrow morning at 08.00

17/06/2011

10.00 Beam coming in 45 minutes

10.45 Starting installation of the Single MaskGEM behind the magnet

13.20 Due to acces problem we lost 40 minutes of beam, door PPE152 must be patrolled by CCC but this takes really too much.

13.30 Beam now is ok thanks to the access problem fixed.

13.32 SPS lost the beam.. we had only few seconds of beam...

14.00 PPE172 in access with key to allow CALICE installation

15:40 CALICE has agreed to let us close area at 17:00, we are on standby until then

17.15 Closing the area, setting up the beam.. some problem.. in contact with CCC...

17.45 CCC finally is saying that they does not know what to do, calling the physics liaison

19.30 No way to get a beam, we have tried all possible things.. at 20.00 Calice is planned to enter and finish the installation.

20.00 Giving acces to CALICE

22:30 We have closed the area and aked for the high intensity beam for H2. 22:30 We have closed the area and aked for the high intensity beam for H2. 22:53 CCC is making an intensity scan for H2 it should take an half of hour. 22:53 CCC is making an intensity scan for H2 it should take an half of hour. 23:37 At the moment we have 500 counts per spill and we are starting data taking. CCC is refilling the LHC now, they promise us to get back to our case afterwards. 23:37 At the moment we have 500 counts per spill and we are starting data taking. CCC is refilling the LHC now, they promise us to get back to our case afterwards. -------

18/06/2011

01:11 We are taking data. The rate is about 1000 events/spill. CCC cannot do more then this. we are taking data on the last HV point.

07:57 we are about to finish point 5 (point 6 done) but the rate is very low : about 1000 events/spill

08.10 Giving access to CALICE for 6 hours, in the mean while around 11.30 we plan to go to H4 upstream and CCC to understand the problem with the oncall liason Ilias 08:00 We stop the datataking due to the accsess planned for 6 hours. All HV is off

9.30 Working at the CCC together with the SPS shifter to retrieve same condition as few days ago

10.30 It seems that the Berillum target was not in place (500mm instead of 300mm), while wobbling settings remained unchanged (even if the working file seems to be set for a very different beam angle), asking access to CALICE just for few minutes to check if with H2B.987 plus_target500mm we can have a nice beam profile

11.30 We succeed in setting a beam with a reasonable rate (actually 40k counts/spill...), H2B.987 works again with the correct target500mm, still some setting is not well understood but once there is some time I am going to play with the conf.

12.00 Beam OFF: CALICE is finishing the installation, let's resume at 15.00 for a very intense data-taking programme together with the magnet!

16.00 Waiting for CALICE to finish the installation, expected in 15 minutes. At 16.30 we plan to have BEAM ON and to be ready for data-taking.

17:30 CALICE left at 4.30 pm, but now we have problems again with the beam. Trying to resolve. Again the rate is 300 Hz/spill.

19.00 Still beam is unstable and with very low rate, we are trying to understand what is wrong but without the experts it is hard

19.45 Access requested by CALICE (25 minutes)

20.05 Starting data taking.. low rate as last night (1k counts/spill).. nothing to do with BEND06, BEND08 and QUAD16..

21.00 Having a look to run 460 (just finished) beam profile is 10cm wide.. like this is difficult to calibrate the singlemaskPrototype installed after the magnet, also from the tracker the beam profile is kind of bended (I mean the beam arrives with a non negligible angle) and then the tracks-reconstruction is poor.. (only 10%...)

21.15 CALICE asked another fast access (10 minutes hopefully..)

21.30 Beam ON.. stopping the HV scan with CF4 to give priority to the "Lorentz angle measurement", now I have got 12k counts / spill, collimators totally open (maybe CALICE is providing us something also..)

22.00 We are just limited by TURBO, it is set to collect data as much as possible.. the plan is to take 500k ev with B=0, and to go on with same statistics with B=0.6,1,1.5,2,2.5,3 up to tomorrow morning at 8.30 when CALICE will come asking an access.

22.30 DAQ crash, error in setting the data on disk, "shifter remember to save data on small (not more than 50k) files, never one huge file!" the TDC prescaled runs very nicely, so timingGEM seems not so bad in terms of alignment.

22.50 We agreed to take one more run for the HV scan (p4) and after take runs with the magnet as planned at 22.00

23.50 We will continue taking data up to tomorrow morning 08.00 when access will be given to CALICE (until 10.00)

19/06/2011

1:30 We noticed, that magnetic field affects PMTs at 0,6T (3~6 counts). We decided to decrease the magnet current to 600A (~0,5T) in order to restore a bit more counts.

2:02 We are continuing data taking at 0,5T (600A), we have around 66 counts per spill.

2:10 No beam, expert called (some power issues). We are waiting for more information.

2:40 Beam is back but the rate is lower (around 40 counts per spill).

3:00 Constant beam problems (interuptions).

3:30 Beam estabilished again.

3:35 Problems with software (several crashes, missing VFAT).

3:55 Problems solved thanks to Andrey, we return to data taking.

6:39 No beam (problem is under investigation), we are waiting for more details.

6:50 PSB problem, we are waiting for more news in 30 min.

7:45 Still no beam, we have collected 7k events over the night. Now, we are switching off the magnet and shutting down the detector's HV.

8:40 No beam, free access to the area. Magnet off (~100A).

12:00pm CALICE needs access till 2:00pm

12.30 We use this access time to mount magnetic shielding to the PMTs, starting installation

13:10 The plan is to go for a Magnetic Scan to see if the shielding is working.

14.00 CALICE asked 2 more hours to finish the installation.

15.40 Moving the single mask up of 1cm to try to improve alignement (from a quick estimation the beam will be bended 1.2cm due to 3T field, radium 166m)

17.15 We are closing the area (late due to CALICE..): goals: check alignement and scintillator rate with magnet

18.30 Magnet test performed. We are doing the plot ScintillatorCount vs. MagneticField, according to the slope, we lose 50% counts at 1.5T.. now ramping down, once possible we enter to adjust alignment according to the runs taken in the mean while.

19.20 Updated plan: move the GE11new back to the middle position, which is less dramatic to align, then in the evening/night we plan to take the 0.6T and 1T.

20.30 Magnet still ramping down.. 800A.. once at low current we plan to enter to put GE11new at middle position and adjust single mask detector for tracking. In the mean while I am playing with beam configuration to get the best rate (pure muon)

21.15 Access together with CALICE (we modify GE11new to the middle position

22.00 End of access, checking alignment and setting the beam

22.10 Taking data for alignment.. and setting the beam, wow now I can see our scaler inside CEDAR, let's scan all the magnet, quadrupole, dipole and choose the best rate.

23.00 CALICE ask another access.. in principle we could check alignment but having a look to the run taken now single mask and GE11new seems to be not so bad (I see a small problem due to some mapping issue but I will solve in few hours..)

23.15 During the access CALICE realized to have some broken piece, they need time for reparation, but now since we have the magnet ready in stanby and the beam configuration ready we are requiring beam.. not simple situation, we are just losing time.

23.45 We have 5000 counts / spill, pure muons!!!! Now we know very well how to set up things here.. but just now the TimingGEM powersupply does not respond anymore, we need an acces.. we postpone data-taking..

24.00 Ok we have fixed the problem, beam ON again, taking another run just to see/check (we confirm that we are just fast!!!!! Actually we have saturated TURBO!)

20/06/2011

00.15 Magnet at 800A (=0.6T), waiting for first "trial fast run" to finish before the high statistics runs..

00.25 Trial run seems ok, we start data-taking with 0.6T, we foresee 500k events (each file 10k), after that we will ramp up to 1T and we will take the same statistics, so that for tomorrow morning 08.00 we should finish.

4:01 Data taking at 0,6T has been finished (0,5 mln events). At the moment we are stabilizing the magnet at 1T.

7:30 We have finished data taking at 1T (0,5 mln events). We are swittching off detectors HV and the magnet.

7:42 Excellent job guys, let's give access now, and meet around 10 am for planning the rest

8.00 CALICE acces, in the mean while we are working on the big scintillator in coincidence behind the magnet

10.30 Plan of today: Data-taking with high statistics without magnet along with TDC (as last night)

10.50 CALICE will finish around 12.00

13.15 Acces finished, setting up data-taking with B=0, we plan 500k events,

13.30 Beam ON, muons 6.5k / spill, which is perfect, data-taking... (but at 15.30 there will be no beam for 1h due to SPS intervention), at 17.00 meeting to decide the plan for the night and for tomorrow.

15.30 No Beam for SPS intervention (1h..)

18.15 Beam is back, restarting data-taking (still B=0 lorentz angle meas.), once finished we plan to rotate the detector.

22.05 the prototype 2 and timing gem have been rotated by 30 degrees wrt the B field and we have again started data taking with B=0 T, hv on P6 (Last point on the efficiency plateau) and high eta..

22.15 Taking calibration run to check the new alignment..

22.40 After few iterations now it seems well aligned.

21/06/2011

00:40 We started data taking B=0, (Proto 2 and Timing rotated by 30 degrees). We wish to collect 0,5 mln events, then we will repeat data taking at B=1,5T.

02:24 rate is ok, data taking in progress, we could switch to B=1.5 at 4.00 and finish data taking at 08.00 if beam is stable. TDC running every hours.

03:55 Data taking has been finished (0,5 mln events B=0), we are ramping up the magnet to 1,5T.

4:30 We are operating at 1,5T. It seems that magnetic field affects us again (low rate ~500 counts per spill).

8:08 We have collected 100k events, we are continuing data taking

10.00 No Beam for a problem with LINAC.. more news soon..?

10.30 Beam back, continue data-taking

13.45 No beam until 14.30

15.00 Beam is back, continuing data-taking

16.30 We had our meeting and decided the plan for today evening/night: we are going to continue the ongoing runs with high statistics at 30degree with 1.5T up to 20.00 this evening. Then we will start an HV scan (choosing corret lat, thr=40, MSPL=3) at 30degree with 1.5T. The scan will be over at midnight, then we will rotate the chamber at 0degree and we will start another HV scan keeping the magnet ON at 1.5T.

21:10 Some problems with the beam "SPS MAINS DOWN: trying to restart", CCC unsure of when we will resume.

21:40 SPS status "Trouble on all beams: Investigating..."

21:45 From LHC Page 1 "No beam from SPS - RF Problem, Expert Working - no time estimate." When the beam does resume we will continue the HV scan at 30 degrees for the next HV point.

22/06/2011

1:03 Still no beam - more news ASAP

1:45 Beam is back, we have started data taking at previous HV point - 5th (High-ETA, 30 degrees rotation).

6:01 HV scan done (15k events per point). We are decreasing magnet's current in order to have safe access to the area.

7:10 Detector rotated (0 deg.), problem with locking the area (xml for point 5 ready).

14:45 We are in access mode as there is no beam and no prognosis when it comes. Plan to continue parasitically to complete the HV scan at 1.5 T

19:54 We could use the beam as CALICE is not running, but we have a problem with the magnet. We (Stefano and Andrey) will see if we can make it work until 10 pm, else we will cancel the late night shifts..stay tuned

21.30 We have added a new trigger. We have scintillator A, scintillator B and u3. We are ramping up the magnet to 2000 Amps(1.5 T). No time measurement can be taken this night. A HV scan is foreseen.

23/06/2011

00.39 recover the previous trigger because we are the main user tonight, it will be a muon beam. HV scan from point 5 to point 1, B=1.5Tesla.

00:56 We have very low rate (~5 counts pers spill).

01:28 We have reloaded MPGD config file in order to have better rate. This did not help a lot (~30 counts per spill). We also called CCC but they can not increase intensity any more, the only thing they promise to do is work on a beam symetry.

07:35 Once we have collected 10k events for point 5 of HV scan , we decided to go to the next points with amount of 1k events - just to have an overview. Otherwise we would be stuck at one of the first xml files of point 4 (rate was the same for whole night ~20 counts per spill).

14:35 CALICE has the acess now

23:05 We have done last three points of HV scan (low statistic 1k events per HV point/15 events per spill).

24/06/2011

We are OFF.. workshop...

25/06/2011

8.00 Muon rate 680counts/spill (on our scintillator we have 70 counts/ spill), the muon rate is anyway totally low, I see still problems in the beam configuration file.

9.00 Run meeting to decide the week-end plan.

10.15 Switching ON detectors, starting long run with high statistics. ( I restored the old configuration file which gives a much more reasonable rate)

27/06/2011

7.30 The long Run started on 25th reached 1,5Million of events!!! And it is still counting..

8.00 CALICE access.. well we take data with lower rate but still >1k counts/spill

9.45 No BEAM (PS problem) 1h minimum

12.00 Beam is not stable.. data-taking is then in stand-by waiting for the beam

16.30 Still no beam.. the program is to finish the missing hv scan, and take a run at fully efficiency with 1.5T to use as reference for the tomorrow's run where we will flip (180deg) all detectors.

17.15 Beam is back.. but target not in and intensity low..

17.16 Beam now is stable, starting latency scan and data-taking

17.30 Beam is not stable.. waiting..

18.00 Problems from the Booster (as all the day).. this time no time estimation, CCC does not know also.. waiting...

21:20 Beam comes and goes ... then again Booster septum problem as before, more news later ...

28/06/2011

02:35 CALICE has an access since few hours (no data taking).

03:19 CALICE has finished but no beam from booster.

04:02 Beam is back 600 counts per spill.

8.00 Data-taking (slow) is continuing.. we expect anyway to stop around 10.00 to flip the detectors, at 10.00 also there will be no BEAM for a while (30min?) to allow some SPS non-planned maintainance.

10.00 Entering to flip the chambers (no beam anyway..)

12.00 Beam is back, detectors re-installed, waiting for the gas to flush (CALICE is continuing having access to solve theirs problems), we plan to resume here at 15.00 and start ramping up..

15.30 Starting ramping up the detectors.. CALICE has a long access.. but we see 800counts / spill.. enough for calibration. We plan start dummy runs at 16.00

16.30 Some delay due to software crash and slow data taking (actually we are taking data along with the CALICE access..)

17.00 OK I had a look to the dummy run "2003", for the alignement (after inverting the detectors by 180degree and it looks ok both ge11new and timingGEM. Now starting 15k events for reference and once finished the plan is to ramp up the magnet up to 1.5T and start a long high stat. run. Eventually if there is time tomorrow we could ramp up to 2T.

22.00 Everything looks fine. We are ramping up the magnet to 1.5 T. We are going to take high statistics in the night.

29.06.2011

00.05 The magnet is now at 1.5 Tesla. We have started taking new data.

2:54 Around 00:30, there was a major gas alarm in the area. The beam has been stoped and the HV switched off, Andrey came on call. After a while fire brigade appeard to investigate the problem. They have found an Isobutane leak in the CALICE setup. They decided to shut down all the gas connections, that goes to the gas point - whole experimental area is without the gas now. Tomorrow the gas expert will come to check the interconnections more in detail. For the safety reasons we also had to switched off the magnet. Once CALICE fix their leak and we get the gas again we will continoue data taking. The report from the fire brigade intervention should be available in the following days. Andrey and Michal.

3:48 Counting room is locked.

7.30 Calling Fire Brigade to restore our gas system..

9:50 Gas is on.

11:52 Magnet on at 1.5Tesla(2000 Amp)

17.00 Finally after several SPS problem we can resume our data-taking.. CALICE has access so rate is really low.. waiting for them to fi

30/06/2011

00:53 I started data taking at ~2T (3000A), ~10 counts per spill.

04:47 Up to now I collected 1k events at 2T. I am switching the magnet back to 1,5T in order to start long run.

05:17 Magnet at 1,5T, starting infinite run (saves data every 10k events).

10.00 Wow we have collected a lots of run..

11:35 Beam dump communication problem causes intermittent beam interlock for north area and CNGS ,so no beam.

14.00 No beam for SPS troubles.. we want to finish the famous hv scan.. still 3 runs.. waiting for the beam..

15.30 Beam OK. Taking data..

18.00 We plan to finish data-taking by this night. The idea is to put the chamber at 30 degree (inverted) and take a long run this night

20.00 hv scan is finished. we are switching off the magnet to be able to rotate the chamber at 30 degree. In the same time ccc has come to rearm the forced door. now it's ok, magnet current is very long to decrease about 1000amps now.

21.40 the chamber has been rotated 30degrees, beam is back and magnet ramping up to 1.5Tesla. the detectors are on high voltage point 6 (most efficency). Start long run data taking in a few minutes.

22.30 Change of program. I have checked the first data and after the rotation we lost alignement.. since we don't want to lose reference alignement we go back to the standard position at zero degree to take long statistics run this night with higher magnetic field. Rotation will be tomorrow.

3:20 I have finished run at 2T (10k events), 0 deg. Ramping up the magnet to 2.5T.

5:27 It's not possible to collect any reasonable data at 2.5T (3333A). The PMTs seem to be affected by the magnetic field - no counts!. I was trying to do the latency scan several times without success (even with Threshold=20 and "1" as a minimum number of events to collect per latency). Now I am ramping the magnet to the highest possible current in which I am still able to see some events.

5:54 Magnet set to highest possible current (3267A, ~2.5T), starting long run.

6:00 No beam, no info available.

6:17 Still no beam, chain 11 tripped, under investigation. Prepared xml file has been uploaded, everything is ready for data taking when the beam appears.

7:02 No beam - expert is coming.

8.00 Beam is back, taking data at 2.5T!!

9.30 No Beam (for 1h..) We stop the DAQ, I have seen some misalignement in the important long run@ 1.5T with inverted chamber, then this is the plan: Magnet ramp down (1h), access to the area and improve alignement, ramp up magnet to 1.5T and re-take the long run.

12.30 Doing some alignement run..

14.00 No beam from PS

15.00 Beam is back: checking the tracker.. something it is not correctly working

16.30 No beam.. vacuum problem

18.15 No beam for the weekend up to next MD.. a bit disappointed, we didn't manage to take again some runs.. but we continue to debug the tracker with the source

09/07/2011

16.30 Beam is arriving..

17.00 Tracker alignement..

18.00 Preparing XML files and starting data taking.

10/07/2011

4:00 4 points of HV scan has been done up to now (4.5kV, 4.4kV, 4.3kV, 4.2kV). We took it for 3 MSPL values (3,2,1), 3 latency values (+1, 0, -1) and 3 threshold values (20, 40, 50). We have had some problems with the noise (~1,2k counts without the beam) and the alignment (we have seen only ~50% of the triggers). The noise problem is partialy solved by plaing with the ground (~90 counts without the beam). I also moved the detector a little bit down to improve the alignment - the result we will see in the following run.

7:58 Currently we are running point 6 (3.9kV) of HV scan. After this run Eraldo wants to do a long one, so we will set the detector back to 4.5 kV and we will take as much statistics as we can using one threshold (50) and one latency (0), while MSPL=3. Once it is finished, we repeat data taking at the rest of the HV points (4.4kV, 4.3kV, 4.2kV, 4.1kV, 4.0kV, 3.9kV, 3.8kV, 3.7kV), also using only one threshold (50) and correct latency (0), while MSPL=3.

Morning: Due to black-out no beam...

16.00 Beam should be back soon (30min?)

19.30 No Beam, experts are working on PS, LINAC, SPS.. we are going to wait few hours more..

NIGHT: No Beam at all..

11/07/2011

07.00 Planning to dismount the ge11new

08.00 The ge11new can stay in the beam area at least for today (anyway we have no beam from the PS..)

12.00 Beam is foreseen in 30 minutes

13.00 Starting data-taking

18.00 Data-taking finished, chamber must be moved out

19.30 Ge11new safely in the control room, ready to be shipped to Meyrin

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatxlsx detailed_runlist.xlsx r1 manage 91.8 K 2011-06-10 - 17:28 StefanoColafranceschi  
Unknown file formatxlsx run_list_CMS2011.xlsx r1 manage 37.0 K 2011-06-10 - 17:29 StefanoColafranceschi  
Microsoft Excel Spreadsheetxls runlist_detailed_v1.xls r1 manage 170.5 K 2011-06-15 - 20:22 StefanoColafranceschi  
Microsoft Excel Spreadsheetxls runlist_detailed_v16.xls r1 manage 407.0 K 2011-07-18 - 07:50 StefanoColafranceschi  
Edit | Attach | Watch | Print version | History: r142 < r141 < r140 < r139 < r138 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r142 - 2011-07-18 - StefanoColafranceschi
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    MPGD All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback